A grating has 10,000 lines per cm. Calculate the maximum number of principal maximum that can be formed for light of wavelength 487 nm.
1
Expert's answer
2020-02-21T09:59:18-0500

Solution. The distance between slits is 


"d=\frac{1cm}{10000}=10^{-6}m"

Using the diffraction grating’s slit formula


"dsin\Theta=m\lambda"

where  m is number of principal maxima; λ=487nm is wavelength; Θ is angle between normal and maximum direction.


"sin\Theta \eqslantless 1"

From the diffraction grating’s formula


"m=\frac{dsin\Theta}{\lambda} \eqslantless \frac{d}{\lambda}=\frac{10^{-6}}{487\times 10^{-9}}=2.053"

Therefore the maximum number of principal maxima is equal to 2.

Answer. 2.